# Globally defined Carroll symmetry of gravitational waves

### v1

@@ Abstract
The locally defined Carroll symmetry of a gravitational wave is extended to a globally defined one. Translations and Carroll boosts associated with two independent globally defined solutions of a Sturm-Liouville equation allow us to describe the motions. The Displacement Memory Effect arises for particular choices of the parameters which yield trajectories with zero momentum. We illustrate our general statements by the Pöschl-Teller profile.
@@ II From Brinkmann to BJR and back
Another remarkable consequence of (II.15) is that when $\det P(u_{0})=0$ for some $u_{0}$, then all transverse trajectories which start from an $x_{0}\in\mathrm{Ker}\,\big(P(u_{0})\big)$ are focused at the origin, and we get a caustic point BoPi89 ; EZHRev . The question will be further discussed in Sec.IV and illustrated in FIG. 2.
@@ III Carroll symmetry in Brinkmann coordinates
The coefficients of ${\partial}_{X}$ and of ${\partial}_{V}$ are indeed the the symmetry generators $\xi$ and $\eta$, mentioned in the Introduction. In Sec.IV they will be spelt out explicitly for Pöschl - Teller .
@@
Both $P$ and $Q$ are solutions of the Sturm-Liouville equation (II.3) EZHRev ; Blau24 as it can be verified by a direct calculation. The Wronskian is $W(P,Q)=P^{\prime}Q-Q^{\prime}P=-1$, therefore $P$ and $Q$ are independent. The expression (III.1) could also be verified independently, by checking the symmetry equation $L_{Y}g=0$ Torre ; GenRelGrav .
@@
IV Illustration by the Pöschl - Teller profile
@@
${\mathcal{A}}\equiv{\mathcal{A}}^{PT}(U)=-\dfrac{4m(m+1)}{\cosh^{2}U}\,.$ (IV.1)
@@
Turning to BJR by (II.4), the Souriau matrix $S$ in (II.7) is regular between two subsequent zeros of $P$ but diverges at the junction points. For $m=1$, for example,
$S_{m=1}(u)=u-\coth u\,$ (IV.2)
@@
$S_{m=2}=\frac{1}{4}\left(u+\frac{3\sinh\left(2u\right)}{2(2-\cosh(2u))}\right)\,$, (IV.3)
@@
Figure 1: The Sturm-Liouville solutions $\bf P$ in (II.3) are DM trajectories written in Brinkmann coordinates with wave numbers $\bf m=1$} and $\bf m=2$}, respectively. The Souriau matrix S diverges where $\bf P=0$, but the pull-back to Brinkmann $\color[rgb]{0,0.5,0}\bf Q=PS$} is regular for all $U$.
A second solution of the Sturm-Liouville equation (II.3), $Q_{m}=P_{m}S_{m}$ , is regular for all $U$ but it is a Legendre function with non-DM behavior. The solutions $P_{m}$ and $Q_{m}$ are independent when the Wronskian does not vanish, $W(P,Q)\neq 0$. For $m=1$ and $2$ we have, for example,
$\begin{array}[]{llclllc}P_{1}(U)&=&{-}\tanh U\,,&\qquad&P_{2}(U)&=&\frac{1}{2}(3\tanh^{2}U-1)\,,\\[6.0pt] Q_{1}(U)&=&{1-U\tanh U},&\qquad&Q_{2}(U)&=&\frac{1}{2}\big(2U-3\tanh U-3U\operatorname{sech}^{2}U\big)\,.\end{array}$ (IV.4)
@@
Coming to symmetries of Pöschl - Teller , the coefficients of the Carroll generators $P$ and $Q$ in (III.1) are shown in FIGs. 3 and 4.
@@
Figure 4: Infinitesmal Carroll generators (IV) with wave number $\bf m=2$}. (a) For a translation, the transverse component ${\bf P}_{2}$ is even, and has two zeros, at $U_{\mp}$. (b) For a boost, ${\bf Q}_{2}={\bf P_{2}S_{2}}$ is odd.
@@
$\Theta_{m=1}^{Brink}$ $=$ $h\frac{\partial}{\partial V}+c\Big({\tanh U}\,\frac{\partial}{\partial X}-(\operatorname{sech}^{2}U)X\frac{\partial}{\partial V}\Big)$ (IV.5)
@@
V Planar representation of the symmetry generators
The symmetry-generating vectors along the trajectories, shown in FIGs.3 and 4, are conveniently viewed in a co-moving tangent space – which is indeed a plane carried along the trajectory. In FIGs. 58 the trajectory is “hidden” in yellow “blobs” and the coordinates $\xi$ and $\eta$ represent the coefficients of ${\partial}_{X}$ and of ${\partial}_{V}$ of the symmetry generators in (III.1), $\Theta^{Brink}={\bm{\xi}}{\partial}_{X}+\eta{\partial}_{V}$, spelt out in (IV.5)-(IV).
@@ V Planar representation of the symmetry generators
Figure 7: Translations act along the $m=2$ DM trajectory by (IV) The trajectory ${\bm{X}}(U)$ with $\bf m=2$} is “hidden” in the yellow blob.
According to FIG.7, the translation generator (with $c=1/2$) is, for $U=$-$\infty$ , $({\bm{\xi}}=1,\,\eta=0)$. FIGs. 3 and 4 confirm that, consistently with (IV.5), the two focal points of the trajectory, at $U_{\mp}$, are left invariant. For $U=+\infty$ the “8-shaped” curve returns to $({\bm{\xi}}=1,\,\eta=0)$, where it started from.
@@
Boosts along the $m=2$ trajectory are shown in. FIG.8. The longest vector on the left (in brown) is, consistently with (IV.5) and (IV) with $b=1/2$ and $c=2$, the boost for $U=$-$\infty$, and the longest one on the right (also in brown) is for $U=\infty$. The two shortest blue arrows (${\bm{\xi}}=\pm 1,{\eta}=0$) show the boost acting as a translations by $\mp b$ at the caustic points $U_{\mp}$.
VI Conclusion
Our results shed further light on the relation of Carroll symmetry and the Memory Effect : the matrix $P$ with initial conditions (II.16) yields our geodesics. The general expression (III.5) which involves both DM and VM is one of our main results.
@@ VI Conclusion
The advantage of the Brinkmann form (III.1), which is one of our principal results, is that it replaces the necessarily singular Souriau matrix $S$ Sou73 by the globally defined Sturm-Liouville solution $Q$. As examplified by the Pöschl - Teller profile, it allowed us to extend the locally given Carrollian Killing vectors to entire spacetime. The clue is that the multiplying by $P$ cures the singularity of the Souriau matrix $S$ and thus that of the BJR coordinate system. The Brinkmann trajectories are focused at the zeros of $\det(P)$, as shown in FIG.2.
The Carroll symmetry generators, first identified by Souriau Sou73 , look more complicated in Brinkmann coordinates as their BJR counterparts do. They are however defined globally. The are directly related to geodesics and allow us to unify the Displacement and Velocity Memory effects by using the two independent solutions $P$ and $Q$ of the Sturm-Liouville problem, see (III.5). The intimate relationship of geodesics with Carroll symmetry is highlighted by their common use of the Sturm-Liouville equation GenRelGrav .
Our results go actually beyond $D=1$ dimensional profiles and the Pöschl - Teller example. Similar results hold for the Scarf profile in $D=2$, which yields a good analytic approximation of flyby ZelPol ; Sila-PLB .
The supersymmetric extension MSUSY sheds further light on the Memory Effect. Further details will be presented in a forthcoming comprehensive review ZEHPR .
@@ Acknowledgements.
The authors are indebted to J. Balog for correspondance. PMZ was partially supported by the National Natural Science Foundation of China (Grant No. 11975320). ME is supported by Istanbul Bilgi University research fund (BAP) with grant no: 2024.01.009 and by The Scientific and Technological Research Council of Turkey (TÜBITAK) under grant number 125F021. PMZ and PAH benefitted of TÜBITAK 2221 grant and hospitality at Bilgi University during their visit in Istanbul.

### v2

@@ Abstract
The locally defined Carroll symmetry of a gravitational wave is extended to a globally defined one by switching to Brinkmann coordinates. Translations and Carroll boosts associated with two independent globally defined solutions of a Sturm-Liouville equation allow us to describe the motions. The Displacement Memory Effect arises for particular choices of the parameters which yield trajectories with zero momentum. The relation to the Schwarzian derivative is highlighted. We illustrate our general statements by the Pöschl-Teller profile.
@@ II From Brinkmann to BJR and back
Another remarkable consequence of (II.15) is that when $\det P(u_{0})=0$ for some $u_{0}$, then all transverse trajectories which start from an $x_{0}\in\mathrm{Ker}\,\big(P(u_{0})\big)$ are focused at the origin, and we get a caustic point BoPi89 ; EZHRev . The question will be further discussed in Sec.V and illustrated in FIG. 2.
@@ III Carroll symmetry in Brinkmann coordinates
The coefficients of ${\partial}_{X}$ and of ${\partial}_{V}$ are indeed the the symmetry generators $\xi$ and $\eta$, mentioned in the Introduction. In Sec.V they will be spelt out explicitly for Pöschl - Teller .
@@
The remarkable fact is that both $P$ and $Q$ are solutions of the Sturm-Liouville equation (II.3) EZHRev ; Jibril ; Blau24 , as it can be verified by a direct calculation. The Wronskian is $W(P,Q)=P^{\prime}Q-Q^{\prime}P=-1$, therefore $P$ and $Q$ are independent. The expression (III.1) could also be verified independently, by checking the symmetry equation $L_{Y}g=0$ Torre ; GenRelGrav .
@@
IV Schwarzian aspects
For simplicity, we work in $D=1$. Our results have an interesting relation with the Schwarzian derivative of a function $f$ of is some real variable $x$ Ovsienko ; Vlahakis ; ZZH ,
${\mathfrak{S}}\big(f(x)\big)=\frac{f^{\prime\prime\prime}(x)}{f^{\prime}(x)}-\frac{3}{2}\left(\frac{f^{\prime\prime}(x)}{f^{\prime}(x)}\right)^{2}\,.$ (IV.1)
We note for later use that the Schwarzian depends only on $f^{\prime}$, not on $f$ itself.
Consider now the Sturm-Liouville equation
$\varphi^{\prime\prime}(x)-{\mathcal{A}}(x)\varphi=0\,$, (IV.2)
cf. (II.3) for both $P$ or for $Q$ collectively. Let then $\varphi_{1}(x)$ and $\varphi_{2}(x)$ be two independent solutions and consider their quotient,
$f(x)=\frac{\varphi_{1}(x)}{\varphi_{2}(x)}\,.$ (IV.3)
It is then shown in Ovsienko that the Sturm-Liouville profile can be recovered from the solutions in terms of the Schwarzian derivative,
${{\mathcal{A}}=-{\scriptstyle{\frac{1}{2}}}\,{\mathfrak{S}}(f)\,.}\;\;$ (IV.4)
Let us now apply these general formulas with the cast $x\leadsto U$, completed by,
$\varphi_{1}(U)=P(U){\quad\text{\small and}\quad}\varphi_{2}(U)=P(U)S(U)\,.$ (IV.5)
Thus
$f=\frac{1}{S}\,.$ (IV.6)
The Wronskian $W(P,Q)$ is a constant, implying that $f^{\prime}$ in (IV.1) is indeed proportional to $S^{-2}.$ Thus by (IV.4) the profile is the Schwarzian derivative of the inverse of the Souriau matrix,
${\mathcal{A}}=-\frac{1}{2}\,{\mathfrak{S}}\left(\frac{1}{S}\right)\,.$ (IV.7)
The form (II.7) of $S$ then implies that $P$ determines also the second solution we called $Q$.
Remarkably, we can also interchange the cast: choosing
$\varphi_{1}(U)=Q=P(U)S(U){\quad\text{\small and}\quad}\varphi_{2}(U)=P(U)\,$ (IV.8)
instead of (IV.5), it is $P$ which drops out, leaving us with the reciprocal relation for
$\tilde{f}=\frac{Q}{P}=S\,$, (IV.9)
instead of $f=P/Q=1/S$ as in (IV.6)  [^4] The singularities of $S$ has will be discussed in the next section for PT. The new cast (IV.8) then yields, once again, a SL eqn. (IV.2) whose profile is now determined by the Souriau matrix  [^5] Eqn. (IV.10) can be viewed as a equation for the Souriau matrix $S$.,
${\mathcal{A}}(U)=-2{\mathfrak{S}}(S)=-2\frac{S^{\prime\prime\prime}(U)}{S^{\prime}(U)}+{3}\left(\frac{S^{\prime\prime}(U)}{S^{\prime(}U)}\right)^{2}\,.$ (IV.10)
Somewhat surprisingly, inserting here $S$ from (II.7), the previous Sturm-Liouville equation (II.3) with the same profile $\mathcal{A}$} is recovered. These results indicate that the interchange
$P,\,Q\to Q,\,P\;\;\;\Rightarrow\;\;\;S\to\frac{1}{S}$ (IV.11)
is a symmetry of the system, as it follows from the reciprocity relation
${\mathfrak{S}}(f)={\mathfrak{S}}\left(\frac{1}{f}\right)$ (IV.12)
of the Schwarzian. The difference between the two casts comes from the initial conditions. Those for $P$ are dictated by the behaviour at $U=$-$\infty$ but those for $Q$ depend also on the integral (II.7).
V Illustration by the Pöschl - Teller profile
@@
${\mathcal{A}}\equiv{\mathcal{A}}^{PT}(U)=-\dfrac{4m(m+1)}{\cosh^{2}U}\,.$ (V.1)
@@
Turning to BJR by (II.4), the Souriau matrix $S$ in (II.7) is regular between two subsequent zeros of $P$, but diverges at the junction points. For $m=1$, for example,
$S_{m=1}(u)=u-\coth u\,$ (V.2)
@@
$S_{m=2}=\frac{1}{4}\left(u+\frac{3\sinh\left(2u\right)}{2(2-\cosh(2u))}\right)\,$, (V.3)
@@
Figure 1: For integer $m$ the Sturm-Liouville solutions $\bf P$ in (II.3) are DM trajectories, written in Brinkmann coordinates with wave numbers $\bf m=1$} and $\bf m=2$}, respectively. The Souriau matrix S diverges where $\bf P=0$, but the pull-back to Brinkmann $\color[rgb]{0,0.5,0}\bf Q=PS$} is regular for all $U$.
$Q_{m}=P_{m}S_{m}$ , a second solution (III.3) of the Sturm-Liouville equation (II.3) shown in FIG.1, is a Legendre function. It is regular for all $U$, vanishes where the Souriau matrix does, and exhibits a VM behaviour where $P_{m}$ is DM.
The solutions $P_{m}$ and $Q_{m}$ are independent when the Wronskian does not vanish, $W(P,Q)\neq 0$. For $m=1$ and $2$ we have, for example,
$\begin{array}[]{llclllc}P_{1}(U)&=&{-}\tanh U\,,&\qquad&P_{2}(U)&=&\frac{1}{2}(3\tanh^{2}U-1)\,,\\[6.0pt] Q_{1}(U)&=&{1-U\tanh U},&\qquad&Q_{2}(U)&=&\frac{1}{2}\big(2U-3\tanh U-3U\operatorname{sech}^{2}U\big)\,.\end{array}$ (V.4)
@@
Coming to symmetries, the coefficients of the Carroll generators $P$ and $Q$ in (III.1) are shown in FIGs. 3 and 4.
@@
Figure 4: Infinitesmal Carroll generators (V) with wave number $\bf m=2$}. (a) For a translation, the transverse component ${\bf P}_{2}$ is even, and has two zeros, at $U_{\mp}$. (b) For a boost, ${\bf Q}_{2}={\bf P_{2}S_{2}}$ is odd.
@@
$\Theta_{m=1}^{Brink}$ $=$ $h\frac{\partial}{\partial V}+c\Big({\tanh U}\,\frac{\partial}{\partial X}-(\operatorname{sech}^{2}U)X\frac{\partial}{\partial V}\Big)$ (V.5)
@@
VI Representation of the symmetry generators
The symmetry-generating vectors along the trajectories, shown in FIGs.3 and 4, are conveniently viewed in a co-moving tangent space – which is a plane carried along the trajectory. In FIGs. 58 the trajectory is “hidden” in yellow “blobs” and the coordinates $\xi$ and $\eta$ represent the coefficients of ${\partial}_{X}$ and of ${\partial}_{V}$ of the symmetry generators in (III.1), $\Theta^{Brink}={\bm{\xi}}{\partial}_{X}+\eta{\partial}_{V}$, spelt out in (V.5)-(V).
@@ VI Representation of the symmetry generators
Figure 7: Translations act along the $m=2$ DM trajectory by (V) The trajectory ${\bm{X}}(U)$ with $\bf m=2$} is “hidden” in the yellow blob.
According to FIG.7, the translation generator (with $c=1/2$) is, for $U=$-$\infty$ , $({\bm{\xi}}=1,\,\eta=0)$. FIGs. 3 and 4 confirm that, consistently with (V.5), the two focal points of the trajectory, at $U_{\mp}$, are left invariant. For $U=+\infty$ the “8-shaped” curve returns to $({\bm{\xi}}=1,\,\eta=0)$, where it started from.
@@
Boosts along the $m=2$ trajectory are shown in. FIG.8. The longest vector on the left (in brown) is, consistently with (V.5) and (V) with $b=1/2$ and $c=2$, the boost for $U=$-$\infty$, and the longest one on the right (also in brown) is for $U=\infty$. The two shortest blue arrows (${\bm{\xi}}=\pm 1,{\eta}=0$) show the boost acting as a translations by $\mp b$ at the caustic points $U_{\mp}$. Combining FIGs. 3 with FIG.5 provides us with the outfolding to 3dimensions, shown in FIG. 9.
Figure 9: The translations implemented along the trajectories with $m=1$ and $m=2$.
The translations implemented along the trajectories vanish where $P(U)=0$. They return to their initial value for $m$ odd and their transverse component changes sign for $m$ even.
Figure 10: Carroll boosts implemented along the trajectories with $m=1$ and $m=2$.
Carroll boosts implemented along the trajectories behave at pure translations at P(U)=0P(U)~=~0. They return to their initial value for $m$ odd and their transverse component changes sign for $m$ even.
Conversely, we duly recover the Pöschl - Teller profile (V.1) from the Sturm-Liouville solutions by using the Schwarzian formulas in sect.IV. We show this here for the simplest, $m=1$ solution (V.4), with the cast
$\varphi_{1}(U)=P(U)=-\tanh U,\qquad\varphi_{2}(U)=Q(U)=1-U\tanh U\,.$ (VI.1)
The Wronskian equals to (1)(-1) everywhere. Then inserting $f^{\prime}(U)=-\cosh^{2}U+1\,$ into the Schwarzian (IV.4) reproduces the $m=1$ profile ${\mathcal{A}}(U)=-\frac{8}{\cosh^{2}(U)}$ in (V.1).
VII Conclusion
Our results shed further light on the intimate relation of Carroll symmetry and the Memory Effect : the matrix $P$ with initial conditions (II.16) yields our geodesics. The general expression (III.5) which involves both DM and VM is one of our main results.
@@ VII Conclusion
The advantage of the Brinkmann form (III.1) is that it replaces the necessarily singular Souriau matrix $S$ Sou73 by the globally defined Sturm-Liouville solution $Q$ – another one of our principal results,
Examplified by the Pöschl - Teller profile, it allowed us to extend the locally given Carrollian Killing vectors to the whole spacetime. The clue is that the multiplying by $P$ cures the singularity of the Souriau matrix $S$ and thus that of the BJR coordinate system. The Brinkmann trajectories are focused at the zeros of $\det(P)$, as shown in FIG.2.
The Carroll symmetry generators, first identified by Souriau Sou73 , look more complicated in Brinkmann coordinates as their BJR counterparts do. They are however defined globally. They are also directly related to geodesics and allow us to unify the Displacement and Velocity Memory effects by making use of the two independent solutions $P$ and $Q$ of the Sturm-Liouville problem, see (III.5). The intimate relationship of geodesics with Carroll symmetry is highlighted by their common use of the Sturm-Liouville equation EZHRev ; GenRelGrav .
Figs. 510 illustrate the curious behavior of the Carroll generators along the trajectories.
Further insight is provided by the Schwarzian framework Ovsienko ; ZZH ; Vlahakis . This approach confirms that $Q=PS$ is indeed a global, independent, 2nd solution of the Sturm-Liouville problem which does not involve the locally defined Souriau matrix $S$.
The supersymmetric extension MSUSY sheds further light on the Memory Effect.
Our results go beyond $D=1$ dimensional profiles and the Pöschl - Teller example. Similar results hold for the Scarf profile in $D=2$, which yields an analytic approximation of flyby ZelPol ; Sila-PLB . Details will be presented in a forthcoming comprehensive review ZEHPR .
@@ Acknowledgements.
The authors are indebted to J. Balog, J. Ben Achour and Z. Silagadze for correspondance. PMZ was partially supported by the National Natural Science Foundation of China (Grant No. 11975320). ME is supported by Istanbul Bilgi University research fund (BAP) with grant no: 2024.01.009 and by The Scientific and Technological Research Council of Turkey (TÜBITAK) under grant number 125F021. PMZ and PAH benefitted of TÜBITAK 2221 grant and hospitality at Bilgi University during their visit in Istanbul.